How much work is done by the system when
2.00 moles of O2 expand from a volume of
3.00 liters to a volume of 6.9 liters against a
constant external pressure of 0.0100 atm?
Answer in units of J
1
Expert's answer
2013-12-16T11:24:34-0500
The system is isobaric, so we can use the following equation for the work in system:
A = pΔV
p = 0.0100 atm - 1013.25 Pa; ΔV = 0.0069 - 0.0030 = 0.0039 m3
A = 1013.25*0.0039 = 3.95 J